OpenStudy (anonymous):

4r-81 factor completely

OpenStudy (id21):

\[4r-81=2^2(\sqrt{r})^2-9^2=(2\sqrt{r}-9)(2\sqrt{r}+9)\]

OpenStudy (id21):

I used formula \[a^2-b^2=(a-b)(a+b)\]

OpenStudy (anonymous):

so thats the whole answerat the end

OpenStudy (id21):

Yes, answer is \[(2\sqrt{r}-9)(2\sqrt{r}+9)\]